Today we were in squeezing mode.
1. Alignment of SC and LO on HD
We checked the alignment of the SC and the LO on EQB1_HD_DIFF_RF_4MHz_mag.
After the alignment with M13, M4 and M6, we reached a magnitude oscillating between 0.00226 and 0.00658, which corresponds to parametric gain of 2.9 (9.25 dB of squeezing producted).
We adjusted the weight of the EQB1_HD_DIFF_AUDIO_RMS_1_3kHz:
ACL_SUM_CH HD_DIFF_AUDIO_RMS_1_3kHz_norm "" 1 273973 HD_DIFF_AUDIO_RMS_1_3kHz
2. Attempt to close the CC fast and coarse on LO_M2 and M4/M6
We worked with the dither line on M4_Z off.
We needed to restart the EQB1_CC process since we were not able to remove a 3Hz fluctuations on the CC corrections. After the restart the oscillation disappeared.
Then we tried to close CC fast and CC coarse but we noticed from the mag signal that the HD was misaligning very fast.
We tried with CC coarse gain -1 and CC gain of 10, 20, 30, 50, 100 but the system was always misaligning.
We tried to send the correction of the CC coarse only on M4 and not M6 but the corrections were drifting very fast.
3. Attempt to close CC coarse loop on FC input mirror
We sent the correction to FCIM_LC_MIR_Z, so the new correction signal is FCIM_LC_MIR_Z_ACT.
We sent a large noise on the input mirror in Z to check the range of the mirror:
NOISE (FCIM_LC_NOISE): sine at 1Hz up to 20 V of amplitude
MIR_Z: oscillation of 19 um (with 20V of noise amplitude)
This is shown in Fig. 1.
To send this noise we used the following function on python (since the other python functions were not working well):
FCIM_LC.sin_noise.AcSinChSet(frequency, amplitude, 0.0, 0.0, 1.0).
4. Diagonalization of MIR_Z, MIR_TX and MIR_TY
We notice that sending a noise on MIR_Z, the spots on FCEB and EQB1_GR cameras was moving a lot in VER and a bit in HOR.
We diagonalized the mirror driving matrix in order to reduce this coupling between different dofs.
We checked the signals MIR_Z, MIR_TX, MIR_TY, changing online via python the weights on the coils of the input mirror, to reduce TX and TY.
To do a final tuning we checked also the oscillation on the cameras X and Y. Even if minimizing TX and TY we had some residual motion on the cameras. This shows that there is a coupling in the optical lever signals. We minimized the beam motions on the cameras with the following driving matrix:
MIR_R = 0.115
MIR_L = 0
MIR_B = 1
MIR_T = -1